The perimeter of a book cover is 14 inches, and the diagonal measures 5 inches. What are the length and width?
Answer by checkley77(12844)

2x+2y=14
2x=14-2y
x=(14-2y)/2
x=7-y
x^2+y^2=5^2
(7-y)^2+y^2=25
49-14y+y^2+y^2=25
2y^2-14y+49-25=0
2y^2-14t+24=0
2(y^2-7y+12)=0
2(y-3)(y-4)=0
y-3=0
y=3 ans. x=4
y-4=0
y=4 ans. x=3
Proof:
2*3+2*4=14
6+8=14
14=14
3^2+4^2=5^2
9+16=25
25=25